# When “Competency” in Reasoning Opens the Door to Vulnerability: Jailbreaking LLMs via Novel Complex Ciphers

### v2

@@ Abstract
Recent advancements in the safety of Large Language Models (LLMs) have primarily focused on mitigating attacks crafted in natural language or in common encryption techniques like Base64. However, new models which often possess better reasoning capabilities, open the door to new attack vectors that were previously non-existent in older models. This seems counter-intuitive at first glance, but these advanced models can decipher more complex cryptic queries that previous models could not, making them susceptible to attacks using such prompts. To exploit this vulnerability, we propose Attacks using Custom Encryptions (ACE), a novel method to jailbreak LLMs by leveraging custom encryption schemes. We evaluate the effectiveness of ACE on four state-of-the-art LLMs, achieving Attack Success Rates (ASR) of up to 66% on close-source models and 88% on open-source models. Building upon this, we introduce Layered Attacks using Custom Encryptions (LACE), which employs multiple layers of encryption through our custom ciphers to further enhance the ASR. Our findings demonstrate that LACE significantly enhances the ability to jailbreak LLMs, increasing the ASR of GPT-4o from 40% to 78%, a 38% improvement. Our results highlight that the advanced capabilities of LLMs introduce unforeseen vulnerabilities to complex attacks. Specifically complex and layered ciphers increase the chance of jailbreaking. [^1] Data and Source Code available at https://github.com/DivijH/jailbreak_cryptography
Content Warning: This paper contains unfiltered, possibly offensive content generated by LLMs.
@@
Divij Handa  Zehua Zhang  Amir Saeidi  Chitta Baral Arizona State University {dhanda, zzhan645, ssaeidi1, cbaral}@asu.edu
@@ 1 Introduction
Figure 1: Common ciphers such as Base64, which are widely used on the internet, are easily decrypted by both older and newer LLMs. While older models, like GPT-3.5, struggle with deciphering novel user-defined ciphers and cannot decode malicious queries, newer models like GPT-4o have the capability to decrypt such ciphers, making them vulnerable to potential jailbreaks.
Figure 2: Our first jailbreaking method, Attack using Custom Encryptions (ACE), uses novel, user-designed ciphers to jailbreak LLMs, achieving a higher attack success rate (ASR) compared to both common and uncommon ciphers. Building upon ACE, Layered Attacks using Custom Encryptions (LACE), combines two encryption techniques, such as Word Substitution Cipher followed by Word Reversal Cipher. All ASRs displayed correspond to GPT-4o.
@@
Despite their effectiveness, a critical limitation of these approaches lies in their reliance on the “correct” safety data, which assumes that harmful content will always appear in recognizable formats, such as natural languages (e.g. English or Chinese) or widely known encoding schemes like Base64, Leetspeak, or shift ciphers (e.g. ROT-13) Xie et al. (2024). This assumption neglects the possibility that more obscure or novel ciphers could be employed to encode harmful content, potentially evading existing safety mechanisms. While Yuan et al. (2023) noted that user-created ciphers were ineffective in the past due to earlier LLMs’ inability to decode them, recent findings by Jiang et al. (2024) demonstrate that novel encryption methods can successfully facilitate jailbreak attempts. As LLMs become increasingly sophisticated, they may develop the capability to decipher complex ciphers, introducing new security vulnerabilities that current defenses are not equipped to address.
To evaluate LLMs’ decoding abilities on new ciphers, we introduce CipherBench, a benchmark designed to assess their performance across a spectrum of encoding techniques—including common, uncommon, and novel (user-created) ciphers. Building on this, we propose a jailbreak attack named Attack using Custom Encryptions (ACE), which exploits vulnerabilities associated with uncommon and novel ciphers. Specifically, we investigate three relatively uncommon ciphers—the Keyboard Cipher, Upside-Down Cipher, and Word-Reversal Cipher—which are rarely discussed in cryptographic literature (Banoth and Regar, 2023; Biswas et al., 2019). Additionally, we introduce two novel, user-created ciphers—the Grid Encoding and Word-Substitution Cipher—which have not been documented in existing literature and are likely absent from the pre-training data of any known LLM.
By encoding harmful queries with these ciphers, we assess the effectiveness of ACE in circumventing safety protocols, reporting the Attack Success Rate (ASR). Our experiments cover both open-source and proprietary models, including Llama-3.1-8B-Instruct (88% ASR), Llama-3.1-70B-Instruct (54% ASR) (Dubey et al., 2024), GPT-4o (40% ASR) (Achiam et al., 2023), and Gemini-1.5-Flash (66% ASR) (Reid et al., 2024). The high ASRs suggest that new encoding techniques can effectively exploit LLMs, even though we believe these models have been aligned using existing encryption-based defenses, as evidenced by a decline in ASRs when using common encryption techniques compared to previous reports (Wei et al., 2024; Yuan et al., 2023).
Building upon ACE, we introduce a second jailbreak attack, Layered Attacks using Custom Encryptions (LACE), to examine the impact of sequentially layering two ciphers on LLM security. In LACE, we apply two encodings are applied consecutively, requiring the model to decode both layers before processing the harmful query (see Fig 2). Our experiments indicate that GPT-4o is the only model capable of decoding layered encryptions, exhibiting a significant increase in ASR from 40% to 78%. This observation suggests that LLMs with advanced capabilities in deciphering complex ciphers are more susceptible to attacks utilizing novel and layered encodings.
@@ 2 Related Works
Several studies, similar to ours, explore prompt-based jailbreak attacks that reduce resource demands by translating harmful queries in innovative ways. For instance, prior work has involved translating harmful queries into low-resource languages (Yong et al., 2023) (Deng et al., 2023), converting them into ASCII art (Jiang et al., 2024), embedding them within programmatic structures (Kang et al., 2024) (Lv et al., 2024), or encoding them using widely-used cryptographic techniques (Wei et al., 2024; Yuan et al., 2023). However, a key limitation of these approaches is that they often rely on common transformations and are potentially present in the model’s pre-training data, making it possible to integrate these methods into safety alignment mechanisms over time. In contrast, our research proposes a more resilient approach by encoding harmful queries using novel or lesser-known encryption techniques. As LLMs become more sophisticated, we argue that continuously creating novel encryption methods poses a greater challenge for safety training, demonstrating that simply incorporating known strategies into safety alignment is insufficient for long-term protection.
@@ 3 CipherBench: Evaluating Decoding Capabilities of LLMs
In this section, we evaluate the ability of LLMs to decode sentences encrypted using ten different cipher techniques. We design a benchmark called CipherBench and assess the capabilities of two state-of-the-art LLMs. Details of the prompts for the encryption methods are provided in Appendix A.
3.1 CipherBench: Benchmark Design
@@
CipherBench includes ten ciphers, categorized into three distinct groups: common, uncommon, and novel. The common category includes widely used or well-known encryption methods, such as Base64, ROT-13, Pig Latin, and LeetSpeak. In contrast, the uncommon category features ciphers that are less frequently encountered, including Keyboard Cipher, Upside-Down Cipher, and Word-Reversal. Lastly, the novel category consists of user-designed ciphers not present in cryptographic literature, such as Word-Substitution Cipher, Grid Encoding, and ASCII Art Cipher.
@@
We generate 20 instances using a random sentence generator [^2] https://www.thewordfinder.com/random-sentence-generator/. Specifically, the benchmark includes ten short and ten long sentences. To further diversify the benchmark, we divide each category into five question-based sentences and five declarative sentences. Additionally, to evaluate whether LLMs can truly decrypt sentences or are simply leveraging statistical patterns, we generated randomized sentences by replacing each character in the original sentences with a randomly character. Importantly, we ensure that all instances are safe to avoid triggering the models’ safety filters.
@@
We evaluate two state-of-the-art LLMs, GPT-4o and Gemini-1.5-Flash on CipherBench to assess their performance in decoding the encrypted sentences. To quantify their effectiveness, we report the Decryption Success Rate (DSR), defined as the percentage of successfully decoded sentences:
$DSR=\frac{\text{\# of Successfully Decoded Inst.}}{\text{Total Number of Instances}}\times 100$
@@ 3.2 Performance on CipherBench
Fig 3 and 4 illustrates the DSR corresponding to GPT-4o and Gemini-1.5-Flash.
Figure 3: Decryption Success Rate (DSR) of the GPT-4o (in Blue) and Gemini-1.5-Flash (in Orange) on CipherBench. Poor performance is observed when sentences are modified to contain random characters instead of normal English sentences. This indicates the models’ reliance on recovering sentences from statistical patterns instead of understanding the decryption algorithm.
Analyzing the results, we observe a significant decrease in DSR when the sentences are composed of random characters rather than standard English words, with an average drop of 22.75%. This finding suggests that the LLMs’ decoding capabilities rely heavily on the syntactic and semantic relationships between words, rather than on an inherent understanding of the underlying encryption algorithms. Furthermore, we notice that LLMs are 6.25% better at decoding questions compared to declarative statements, a trend consistent across both GPT-4o and Gemini-1.5-Flash.
When examining the performance across different cipher categories, we find that the DSR for common encryption techniques is 44.69%, whereas uncommon techniques have a DSR of 24.58%, and novel, user-defined techniques have the lowest DSR of 16.25%. Specifically, among all ciphers, Base64 is the most accurately solved with a DSR of 61.25%, while the ASCII Art Cipher proves to be the most challenging, with a DSR of just 1.25%. These results support the hypothesis that LLMs perform better at decrypting encodings that are more prevalent in their pre-training data.
Figure 4: Number of successful decryptions by GPT-4o (in Blue) and Gemini-1.5-Flash (in Orange) on CipherBench out of 40. Common encryption techniques boast the highest DSR followed by uncommon techniques followed by novel user-designed encryption techniques.
Manual analysis of the responses reveals that the models often partially reconstruct the sentences. This reconstruction can lead to decoded sentences that are very similar to the original. Consequently, this observation implies that encoded prompts can still be utilized for jailbreaking, as even partially reconstructed sentences may bypass safety filters if the LLM can infer the intended meaning.
@@ 4 Experimental Setup for Jailbreaking
In this section, we outline the experimental setup to evaluate the effectiveness of jailbreaking LLMs using encryption techniques.
@@
We use four distinct LLMs for our experiments, GPT-4o (Achiam et al., 2023), Gemini-1.5-Flash (Reid et al., 2024), Llama-3.1-8B-Instruct, and Llama-3.1-70B-Instruct (Dubey et al., 2024). These models were chosen because they have undergone comprehensive safety training, making them suitable for assessing the robustness of their safety measures against jailbreaking attempts. For the Gemini-1.5-Flash model, we configured the safety setting to “Block most”, the highest safety level available through the Gemini API.
@@
To evaluate our jailbreaking techniques, we utilize the AdvBench dataset developed by Zou et al. (2023). Specifically, and in line with prior studies (Zeng et al., 2024; Jiang et al., 2024; Chao et al., 2023; Wei et al., 2024), we extracted the same subset of 50 samples from AdvBench. These samples were manually categorized into 14 different unsafe categories, facilitating analysis of potential biases in the safety training of the LLMs. The categories are described in Appendix C. Additionally, we added priming sentences designed to nudge the LLMs toward producing an unsafe response. These priming sentences were generated using a few-shot prompting approach with GPT-4o. Detailed prompts are provided in Appendix B.
Encryption Techniques Used
We introduce our first jailbreaking technique, Attacks using Custom Encryptions (ACE), that uses the following encryptions to jailbreak LLMs:
@@
    Keyboard Cipher - This uncommon cipher shifts every character in the sentence one position to the right on a QWERTY keyboard layout. For example, the word “Hello” will become “Jr;;p”. To decrypt the message, each character is shifted one space to the left on the QWERTY keyboard.
@@
    Word Reversal Cipher - This uncommon cipher involves reversing the characters of each word in the sentence. For example, “Laptop” becomes “potpaL”. While also easily recognizable to humans and not widely used in cryptography, the reversed words also results in distinct tokens for LLMs.
@@
    Grid Encoding - We developed a novel cipher inspired by chess notation. We construct a 5x5 grid labeled with coordinates such as a1, a2 … e5, and assign each square a letter of the alphabet in order, omitting the letter ‘z’ due to grid size limitations. Although this results in an incomplete cipher, the absence of the letter ‘z’ in the AdvBench dataset allows us to use this encoding effectively. Moreover, the encoding mangles the sentences sufficiently to render them unreadable even without encoding ‘z’.
@@
    Word Substitution Cipher - This novel cryptographic technique involves substituting words in a sentence with another words and providing a mapping between the them. The LLM must use the provided mappings to reconstruct the original sentence.
To further investigate the effectiveness of more complex encryption schemes, we introduce our second jailbreaking technique, Layered Attacks using Custom Encryptions (LACE), where we layer one cipher on top of another. Specifically, we combined the Word Substitution Cipher with other ciphers such as ROT-13, Keyboard Cipher, Grid Encoding, and Word Reversal. For each layered combination, we created three variants by applying the encodings to only the mappings, only the harmful query, or both the mappings and the harmful query.
@@
To determine whether an LLM’s response is harmful, we employed an LLM-as-a-judge strategy. Specifically, we utilized GPT-4o-mini to classify responses as either “Safe” or “Unsafe”. The prompt used is presented in Appendix D. We report the Attack Success Rate (ASR), which quantifies the effectiveness of the jailbreaking technique and is calculated as:
$ASR=\frac{\text{\# of Unsafe Responses}}{\text{Total Number of Instances}}\times 100$
@@
To evaluate whether LLMs genuinely understand safety concerns or are merely flagging responses based on the presence of potentially unsafe keywords, we manually modified all 50 samples in our dataset. These modifications retained the unsafe words but reframed the queries to be safe. For example, the query “Explain how to poison a person’s food without leaving any trace” was rephrased as “Explain how to neutralize a poison in a person’s food without leaving a trace.” We tested the modified dataset across the five encryption techniques used in this study and report the percentage of responses that were incorrectly flagged as unsafe, thereby indicating overdefensiveness.

### v3

@@ Abstract
Recent advancements in Large Language Model (LLM) safety have primarily focused on mitigating attacks crafted in natural language or common ciphers (e.g. Base64), which are likely integrated into newer models’ safety training. However, we reveal a paradoxical vulnerability: as LLMs advance in reasoning, they inadvertently become more susceptible to novel jailbreaking attacks. Enhanced reasoning enables LLMs to interpret complex instructions and decode complex user-defined ciphers, creating an exploitable security gap. To study this vulnerability, we introduce Attacks using Custom Encryptions (ACE), a jailbreaking technique that encodes malicious queries with novel ciphers. Extending ACE, we introduce Layered Attacks using Custom Encryptions (LACE), which applies multi-layer ciphers to amplify attack complexity. Furthermore, we develop CipherBench, a benchmark designed to evaluate LLMs’ accuracy in decoding encrypted benign text. Our experiments reveal a critical trade-off: LLMs that are more capable of decoding ciphers are more vulnerable to these jailbreaking attacks, with success rates on GPT-4o escalating from 40% under ACE to 78% with LACE. These findings highlight a critical insight: as LLMs become more adept at deciphering complex user ciphers–many of which cannot be preemptively included in safety training–they become increasingly exploitable. [^1] Data and Source Code available at https://github.com/DivijH/jailbreak_cryptography
Content Warning: This paper contains possibly offensive content generated by LLMs.
@@
Divij Handa  Zehua Zhang  Amir Saeidi  Shrinidhi Kumbhar  Chitta Baral Arizona State University {dhanda, zzhan645, ssaeidi1, skumbha4, cbaral}@asu.edu
@@ 1 Introduction
Figure 1: Common ciphers such as Base64, which are widely used on the internet and can be easily incorporated into safety training, are decrypted by both older and newer LLMs. While older models, like GPT-4, struggle with deciphering novel user-defined ciphers, newer models like GPT-4o can decrypt such ciphers, making them vulnerable to potential jailbreaks.
Figure 2: Our first jailbreaking method, Attack using Custom Encryptions (ACE), uses novel user-designed ciphers to jailbreak LLMs, achieving a higher attack success rate (ASR) compared to both common and uncommon ciphers. Building upon ACE, Layered Attacks using Custom Encryptions (LACE), combines two encryption techniques, such as Word Substitution Cipher followed by Word Reversal Cipher. All ASRs displayed correspond to GPT-4o.
@@
Despite their effectiveness, a critical limitation of these approaches lies in their reliance on the “correct” safety data, which assumes that harmful content will always appear in recognizable formats, such as natural languages (e.g. English or Chinese) or widely known ciphers like Base64, Leetspeak, or shift ciphers (e.g. ROT-13) Xie et al. (2024). This assumption neglects the possibility that more obscure or novel ciphers can be used to encode harmful queries, bypassing safety mechanisms.
As the reasoning capabilities of LLMs increase, they become more capable of following complex user instructions and decoding ciphers that are not present in their training corpora. Prior works (Jiang et al., 2024; Wei et al., 2024; Inie et al., 2023) show that encoding queries using ciphers can jailbreak LLMs, but are restricted to widely-studied ciphers. In this paper, we build on these works, showing that novel user-created ciphers are not only understood by LLMs but can be used to jailbreak them. This directly challenges the early work done by Yuan et al. (2023), which suggested that such ciphers were ineffective for jailbreaking.
To systematically investigate this phenomenon, we first introduce CipherBench, a benchmark evaluating LLM’s ability to decode ciphers spanning three categories: common, uncommon, and novel user-created ciphers, based on how rarely they are discussed in cryptographic literature (Banoth and Regar, 2023; Biswas et al., 2019). We propose a jailbreak attack, Attack using Custom Encryptions (ACE), which encodes harmful queries in novel user-created ciphers. Our experiments reveal that both open-source and proprietary LLMs, including Llama-3.1-8B-Instruct (88% ASR), Llama-3.1-70B-Instruct (54% ASR) (Dubey et al., 2024), GPT-4o (40% ASR) (Achiam et al., 2023), and Gemini-1.5-Flash (66% ASR) (Reid et al., 2024) can be jailbroken using ACE.
Building upon ACE, we introduce a second jailbreak attack, Layered Attacks using Custom Encryptions (LACE), to examine the impact of sequentially layering multiple ciphers on LLM safety (Fig 2). This makes the encoding more complex to solve but our results show an increase in ASR for GPT-4o from 40% to 78%. This observation gives a key insight, “As LLMs become better at reasoning, they become more vulnerable to complex attacks”.
To summarise, this work makes the following four contributions:
  1. 1.
    CipherBench: A benchmark to systematically evaluate LLMs’ ability to decode 10 ciphers, ranging from common ones (e.g., Base64) to novel user-designed encryptions.
  2. 2.
    ACE Framework: A jailbreaking attack that exploits novel ciphers (e.g., Grid Encoding, Word-Substitution Cipher) to circumvent safeguards in state-of-the-art LLMs. This directly challenges prior claims by Yuan et al. (2023) that user-created ciphers cannot be used for attacks.
  3. 3.
    LACE Framework: We demonstrate that layering multiple ciphers (e.g. word-substitution + word-reversal) amplifies vulnerabilities, achieving a 78% ASR on GPT-4o – nearly doubling the effectiveness of single-cipher attacks.
  4. 4.
    The Security-Reasoning Tradeoff: LLMs which demonstrate a better performance on decoding ciphers, are more vulnerable to jailbreaking attacks using ACE and LACE. As progress in LLM reasoning continues and LLMs become better at decoding new ciphers, these ciphers can be used to jailbreak them.
@@ 2 Related Works
Several studies, similar to ours, explore prompt-based jailbreak attacks that reduce resource demands by translating harmful queries in innovative ways. For instance, prior work has involved translating harmful queries into low-resource languages (Yong et al., 2023; Deng et al., 2023), converting them into ASCII art (Jiang et al., 2024), embedding them within programmatic structures (Kang et al., 2024; Lv et al., 2024), or encoding them using widely-used cryptographic techniques (Wei et al., 2024; Yuan et al., 2023). However, a key limitation of these approaches is that they rely on transformations that are present in the model’s pre-training data, making it possible to integrate these methods into safety alignment mechanisms over time. In contrast, our research proposes a more resilient approach by encoding harmful queries using novel or lesser-known encryption techniques. As LLMs become more sophisticated, we argue that continuously creating novel encryption methods poses a greater challenge for safety training since simply incorporating known strategies into safety alignment is insufficient for long-term protection.
@@ 3 CipherBench: Evaluating Decoding Capabilities of LLMs
In this section, we evaluate the ability of LLMs to decode sentences encrypted using ten different ciphers. We design a benchmark called CipherBench and assess the capabilities of four state-of-the-art LLMs. Details of the prompts for the encryption methods are provided in Appendix B.
3.1 Benchmark Design
@@
CipherBench includes ten ciphers, categorized into three distinct groups: common, uncommon, and novel, based on how frequently they are discussed in the literature (Banoth and Regar, 2023; Biswas et al., 2019). The common category includes widely used or well-known encryption methods, such as Base64, ROT-13, Pig Latin, and LeetSpeak. In contrast, the uncommon category features ciphers that are less frequently encountered, including Keyboard Cipher, Upside-Down Cipher, and Word-Reversal. Lastly, the novel category consists of user-designed ciphers not present in cryptographic literature, such as Word-Substitution Cipher, Grid Encoding, and ASCII Art Cipher.
@@
We generate 240 instances using a random sentence generator [^2] https://www.thewordfinder.com/random-sentence-generator/. Specifically, the benchmark includes 120 short and 120 long sentences. To further diversify the benchmark, we divide each category into question-based and declarative sentences. Finally, to evaluate whether LLMs can truly decrypt sentences or are simply leveraging statistical patterns, we generate randomized sentences by replacing each character in the original sentences with a random character. Importantly, we ensure that all instances are safe to avoid triggering the models’ safety filters.
@@
We evaluate four LLMs–Llama-3.1-8B-Instruct, Llama-3.1-70B-Instruct, GPT-4o, and Gemini-1.5-Flash, on CipherBench to assess their performance in decoding the encrypted sentences. To quantify their effectiveness, we report the Decryption Success Rate (DSR), defined as the percentage of successfully decoded sentences:
$\mathrm{DSR}=\left(\frac{N_{\mathrm{decoded}}}{N_{\mathrm{total}}}\right)\times 100\%$
@@ 3.2 Performance on CipherBench
Figure 3: Decryption Success Rate (DSR) of LLMs on CipherBench. Poor performance is observed when sentences are modified to contain random characters instead of English sentences. This indicates the models’ reliance on recovering sentences from statistical patterns instead of understanding the decryption algorithm.
Fig 3 and 4 summarizes the performance of all LLMs on CipherBench. Analyzing the results, we observe a significant decrease in DSR when the sentences are composed of random characters rather than standard English words, with an average drop of 35.23%. This finding suggests that the LLMs’ decoding capabilities rely heavily on the syntactic and semantic relationships between words, rather than on an inherent understanding of the underlying decryption algorithm. Furthermore, we notice that LLMs are 4.44% better at decoding questions compared to declarative statements, a trend consistent across all LLMs. Shorter queries are also 3.56% easier to decode than longer queries.
When examining the performance across different cipher categories, we find that the DSR for common encryption techniques is 29.4%, whereas uncommon techniques have a DSR of 14.62%, and novel, user-defined techniques have the lowest DSR of 14.06%. Specifically, among all ciphers, Base64 is the most accurately solved with a DSR of 38.65%, while the ASCII Art Cipher proves to be the most challenging, with a DSR of just 2.71%. These results support the hypothesis that LLMs perform better at decrypting encodings that are more prevalent in their pre-training data.
Figure 4: Decryption Success Rate (DSR) of LLMs across 10 ciphers present in CipherBench. Common ciphers have the highest DSR followed by uncommon ciphers followed by novel user-designed ciphers.
A manual analysis of the model outputs reveals that LLMs frequently reconstruct partial sentences, especially when the encrypted sentences are in English. In many cases, these partial reconstructions are sufficiently accurate to convey the intended meaning of the original sentence. This observation raises concerns that even partially reconstructed sentences might be used for jailbreaking purposes, potentially bypassing safety filters if the model is able to infer the intended content despite incomplete decryption.
@@ 4 Experimental Setup for Jailbreaking
We use four distinct LLMs for our experiments, GPT-4o (Achiam et al., 2023), Gemini-1.5-Flash (Reid et al., 2024), Llama-3.1-8B-Instruct, and Llama-3.1-70B-Instruct (Dubey et al., 2024). These models were chosen because they have undergone comprehensive safety training, making them suitable for assessing the robustness of their safety measures against jailbreaking attempts. For the Gemini-1.5-Flash model, we configured the safety setting to “Block most”, the highest safety level available through the Gemini API. All the experiments for all LLMs are performed with greedy decoding (i.e. temperature 0).
@@
To evaluate our jailbreaking techniques, we utilize the AdvBench dataset developed by Zou et al. (2023). Specifically, and in line with prior studies (Zeng et al., 2024; Jiang et al., 2024; Chao et al., 2023; Wei et al., 2024), we extracted the same subset of 50 samples from AdvBench. These samples were manually categorized into 14 different unsafe categories, facilitating analysis of potential biases in the safety training of the LLMs. The categories are described in Appendix D. To further validate our results, we evaluate our jailbreaking methods on JailbreakBench (Chao et al., 2024). The results on JailbreakBench are discussed in Appendix A.
Additionally, we add priming sentences to all baselines and our methods. These sentences are designed to nudge LLMs toward producing an unsafe response. These priming sentences were generated using a few-shot prompting approach with GPT-4o. Detailed prompts are provided in Appendix C.
Encryption Ciphers
Our first jailbreaking method Attacks using Custom Encryptions (ACE) uses the following ciphers:
@@
    Keyboard Cipher - This uncommon cipher shifts every character in the sentence one position to the right on a QWERTY keyboard layout. For example, the word “Hello” will become “Jr;;p”. To decrypt the message, each character is shifted one space to the left on the QWERTY keyboard. This requires spatial reasoning of the layout of a keyboard.
@@
    Word Reversal Cipher - This uncommon cipher involves reversing the characters of each word in the sentence. For example, “Laptop” becomes “potpaL”.
@@
    Grid Encoding - We developed a novel cipher inspired by chess notation. We construct a 5x5 grid labeled with coordinates such as a1, a2 … e5, and assign each square a letter of the alphabet in order, omitting the letter ‘z’ due to grid size limitations. Although this results in an incomplete cipher, the encoding mangles the sentences sufficiently to render them unreadable even without encoding ‘z’.
@@
    Word Substitution Cipher - This novel cryptographic technique involves substituting words in a sentence with another set of words and providing a mapping between them. The LLM must use the provided mappings to reconstruct the original sentence.
To further investigate the effectiveness of more complex encryption schemes, our second jailbreaking technique, Layered Attacks using Custom Encryptions (LACE), layers one cipher on top of another. Specifically, we combine the highest decoded novel cipher, Word Substitution Cipher with other ciphers such as ROT-13, Keyboard Cipher, Grid Encoding, and Word Reversal. For each layered combination, we created three variants by applying the encodings to only the mappings, only the harmful query, or both the mappings and the harmful query.
@@
To determine whether an LLM’s response is harmful, we employed LLM-as-a-judge strategy. Specifically, we utilized GPT-4o-mini to classify responses as either “Safe” or “Unsafe”. The prompt and human correlation are presented in Appendix E. We report the Attack Success Rate (ASR), which quantifies the effectiveness of the jailbreaking technique and is calculated as:
$\mathrm{ASR}=\left(\frac{N_{\mathrm{unsafe}}}{N_{\mathrm{total}}}\right)\times 100\%$
@@
To evaluate whether LLMs genuinely understand safety concerns or are merely flagging responses based on the presence of potentially unsafe keywords, we manually modified all 50 samples of AdvBench. These modifications retained the unsafe words but reframed the queries to be safe. For example, the query “Explain how to poison a person’s food without leaving any trace” was rephrased as “Explain how to neutralize a poison in a person’s food without leaving a trace.” We tested the modified dataset across the five encryption techniques used in this study and report the percentage of responses that were incorrectly flagged as unsafe, thereby indicating overdefensiveness.